OpenStudy (anonymous):

**MEDAL** I'm creating flashcards to help me study for the ASVAB. Please help me with figuring out the correct formulas! The area of one circle is four times as large as a smaller circle with a radius of 3 inches. The radius of the larger circle is A. 12 inches. B. 9 inches. C. 8 inches. D. 6 inches. <-- Formula?

OpenStudy (hang254):

area of circle = pi(r^2)

OpenStudy (anonymous):

@hang254 = 3.14(3^2) right?

OpenStudy (hang254):

correct

OpenStudy (hang254):

just remember PEMDAS

OpenStudy (anonymous):

@hang254 So with using PEMDAS: (3^2) = 9 3.14(9) = 28.2 ...ahh, what did I do wrong?

OpenStudy (hang254):

its asking for the radius of the circle 4 times larger, so you need to find the area of that circle using the given radii, then divide by 4

OpenStudy (hang254):

it should equal 28.2 after dividing by 4

OpenStudy (hang254):

thats how you know its 4 times larger

OpenStudy (anonymous):

@hang254 Soo...you're saying that 4(9pi) = 36pi